# Taxonomy and Estimation of Multiple Breakpoints in High-Dimensional Factor Models

### v2

@@ 5 Simulation
Table 3: Simulated root mean squared errors (RMSEs) and mean absolute errors (MAEs) of $\hat{k}_{QML}$, $\hat{k}_{BKW}$ and $\hat{k}_{MS}$, under DGP 1.C.
$N,T$ $(\hat{k}_{1},\hat{k}_{2})_{QML}$ $(\hat{k}_{1},\hat{k}_{2})_{BKW}$ $(\hat{k}_{1},\hat{k}_{2})_{MS}$
RMSE MAE RMSE MAE RMSE MAE
$\rho=0$ $\alpha=0$ $\beta=0$
100,100 (1.016,0.336) (0.473,0.097) (11.441,9.371) (7.917,3.235) (17.860,5.496) (12.216,2.775)
100,300 (0.915,0.302) (0.436,0.077) (12.283,1.804) (6.219,0.528) (33.254,15.893) (24.851,4.140)
300,300 (0.623,0.152) (0.238,0.023) (10.671,1.417) (5.222,0.425) (42.555,4.143) (28.226,0.878)
300,600 (0.525,0.176) (0.202,0.029) (7.939,0.807) (3.896,0.319) (69.553,5.166) (52.161,1.666)
600,600 (0.428,0.084) (0.147,0.007) (8.311,0.773) (4.289,0.289) (80.610,0.244) (65.299,0.030)
$\rho=0.7$ $\alpha=0$ $\beta=0$
100,100 (0.932,0.272) (0.373,0.058) (18.067,21.688) (14.712,14.044) (14.375,5.354) (9.513,2.835)
100,300 (0.925,0.207) (0.355,0.035) (33.006,21.240) (20.774,6.111) (31.273,25.690) (23.129,9.708)
300,300 (0.519,0.100) (0.179,0.010) (33.220,20.436) (20.490,5.413) (35.426,3.043) (25.134, 0.769)
300,600 (0.503,0.105) (0.171,0.011) (33.781,4.921) (17.369,1.024) (71.489,1.827) (53.703,1.585)
600,600 (0.358,0.105) (0.100,0.009) (34.375,7.976) (16.771,1.160) (96.833,0.316) (83.300,0.100)
$\rho=0$ $\alpha=0.3$ $\beta=0$
100,100 (1.146,0.358) (0.548,0.104) (12.478,8.471) (8.224,3.000) (19.026,5.281) (13.401,2.918)
100,300 (1.019,0.310) (0.449,0.080) (10.994,2.340) (5.665,0.500) (32.237,17.604) (25.454,4.403)
300,300 (0.584,0.182) (0.239,0.031) (12.211,2.303) (5.772,0.503) (34.548,1.604) (25.883,1.120)
300,600 (0.589,0.182) (0.237,0.031) (8.609,0.755) (4.184,0.288) (59.920,1.953) (44.780,1.770)
600,600 (0.449,0.095) (0.152,0.009) (8.119,0.898) (3.966,0.333) (61.981,2.037) (47.405,1.917)
$\rho=0$ $\alpha=0$ $\beta=0.3$
100,100 (0.996,0.300) (0.452,0.080) (12.469,7.292) (8.288,2.384) (15.848,4.817) (10.581,2.675)
100,300 (0.998,0.305) (0.449,0.085) (9.940,1.243) (5.155,0.475) (32.588,20.696) (25.027,6.211)
300,300 (0.621,0.190) (0.229,0.034) (11.364,1.101) (5.958,0.438) (38.026,4.607) (22.914,1.177)
300,600 (0.550,0.170) (0.224,0.027) (8.689,0.784) (4.202,0.298) (71.898,2.406) (53.062,1.562)
600,600 (0.458,0.114) (0.160,0.013) (8.849,0.912) (4.247,0.354) (70.002,0.633) (55.300,0.200)
$\rho=0.7$ $\alpha=0.3$ $\beta=0.3$
100,100 (1.097,0.249) (0.467,0.054) (17.995,21.305) (14.608,13.565) (14.184,3.582) (9.871,1.809)
100,300 (0.982,0.278) (0.412,0.053) (35.058,24.497) (22.097,7.168) (32.378,24.107) (25.337,9.121)
300,300 (0.602,0.138) (0.198,0.019) (31.680,20.713) (19.468,5.404) (33.039,2.012) (25.856,1.414)
300,600 (0.518,0.127) (0.180,0.016) (35.115,8.706) (17.147,1.419) (53.368,1.914) (38.673,1.619)
600,600 (0.486,0.123) (0.150,0.013) (31.109,9.982) (15.349,1.221) (59.029,2.181) (45.495,1.841)
DGP 1.D Let $\Lambda_{1}=\Lambda$, $\Lambda_{2}=2\Lambda$ and $\Lambda_{3}=\Lambda$. In this case, the number of factors remains constant across regimes, and the changes in factor loading are purely proportional, corresponding to a setting with two Type 2 breaks. Table 4 presents the performance of the competing estimators. The table clearly shows that serial correlation in factors significantly impacts the BKW method. When $\rho=0.7$, the RMSEs and MAEs of $\hat{k}_{BKW}$ are substantially larger compared to when $\rho=0$. The MS method also fails to produce accurate estimates in this case, as both its RMSEs and MAEs gradually increase with sample size. In contrast, our QML estimator exhibits much smaller RMSE and MAE. These simulation results support our Theorem 2 (ii), which claims that the estimation error of the QML estimator is $O_{p}(1)$ under rotational change. While $\hat{k}_{BKW}$ also has $O_{p}(1)$ estimation errors, the QML estimator clearly outperforms it based on these simulation results.
Table 4: Simulated root mean squared errors (RMSEs) and mean absolute errors (MAEs) of $\hat{k}_{QML}$, $\hat{k}_{BKW}$ and $\hat{k}_{MS}$, under DGP 1.D.
$N,T$ $(\hat{k}_{1},\hat{k}_{2})_{QML}$ $(\hat{k}_{1},\hat{k}_{2})_{BKW}$ $(\hat{k}_{1},\hat{k}_{2})_{MS}$
RMSE MAE RMSE MAE RMSE MAE
$\rho=0$ $\alpha=0$ $\beta=0$
100,100 (3.496,3.109) (1.575,1.619) (10.317,10.069) (6.108,5.930) (20.658,20.860) (17.500,16.750)
100,300 (2.054,2.109) (1.159,1.142) (6.035,5.698) (2.851,2.448) (47.356,42.423) (37.796,35.259)
300,300 (2.215,2.111) (1.184,1.173) (6.115,5.521) (2.805,2.475) (47.817,41.938) (38.664,34.751)
300,600 (1.888,1.859) (1.859,1.006) (4.273,4.993) (2.078,2.368) (94.145,95.593) (76.831,77.552)
600,600 (2.165,1.987) (1.063,1.017) (5.295,3.856) (2.442,2.000) (95.014,86.729) (77.629,71.579)
$\rho=0.7$ $\alpha=0$ $\beta=0$
100,100 (8.479,8.603) (4.521,4.592) (15.649,16.261) (11.746,12.495) (18.690,18.562) (15.064,15.018)
100,300 (5.341,6.392) (2.765,3.037) (28.486,29.621) (15.658,16.752) (46.643,45.121) (38.320,37.091)
300,300 (6.208,5.149) (3.050,2.782) (29.165,27.955) (16.648,15.486) (46.087,46.810) (37.729,38.957)
300,600 (5.058,5.173) (2.553,2.670) (22.849,23.002) (8.809,9.277) (91.392,93.097) (71.639,75.528)
600,600 (5.695,5.415) (2.832,2.716) (20.543,19.428) (8.459,7.793) (88.560,92.720) (68.569,75.343)
$\rho=0$ $\alpha=0.3$ $\beta=0$
100,100 (3.529,3.143) (1.612,1.657) (10.447,10.092) (6.234,5.912) (19.241,17.709) (16.067,13.640)
100,300 (2.057,2.030) (1.145,1.093) (5.867,5.986) (2.715,2.528) (47.429,44.239) (37.707,36.320)
300,300 (2.225,2.114) (1.192,1.167) (6.159,5.722) (2.805,2.554) (47.924,44.531) (38.340,36.859)
300,600 (1.938,1.895) (1.026,1.027) (4.256,4.964) (2.126,2.354) (96.138,96.734) (78.785,80.658)
600,600 (1.846,1.993) (0.987,1.055) (4.536,4.712) (2.252,2.201) (98.905,94.103) (82.469,76.454)
$\rho=0$ $\alpha=0$ $\beta=0.3$
100,100 (3.358,3.157) (1.486,1.667) (10.340,10.413) (6.117,6.219) (21.086,19.073) (18.085,15.451)
100,300 (3.365,3.380) (1.560,1.641) (10.121,9.705) (6.014,5.840) (20.953,17.881) (18.039,15.115)
300,300 (2.148,2.093) (1.170,1.163) (6.112,5.441) (2.796,2.466) (48.737,42.597) (39.724,35.161)
300,600 (1.912,1.882) (1.003,1.009) (4.229,4.924) (2.081,2.317) (96.557,92.987) (78.411,76.160)
600,600 (1.924,2.038) (1.012,1.086) (4.574,4.716) (2.276,2.222) (97.734,91.644) (81.787,74.427)
$\rho=0.7$ $\alpha=0.3$ $\beta=0.3$
100,100 (8.565,8.408) (4.572,4.490) (15.741,16.328) (11.771,12.532) (16.173,16.614) (13.441,13.643)
100,300 (5.619,6.389) (2.914,3.094) (28.825,30.149) (16.010,17.198) (45.309,44.115) (36.507,36.493)
300,300 (6.326,5.310) (3.101,2.892) (29.341,28.247) (16.790,15.777) (41.986,44.212) (33.615,36.755)
300,600 (4.921,5.171) (2.499,2.685) (22.760,22.752) (8.788,9.281) (93.980,84.635) (74.508,67.463)
600,600 (5.691,5.362) (2.836,2.683) (19.944,19.463) (8.339,7.729) (82.489,86.675) (66.011,70.527)
DGP 1.E In this setup, we consider a case involving Type 2 changes with singular $B_{j}$ matrices. The loadings $\lambda_{i}$ are i.i.d. $N(0,\frac{1}{r}I_{r})$ across $i$, and $\Lambda=(\lambda_{1},\cdots,\lambda_{N})^{{}^{\prime}}$. For the first regime, we set $\Lambda_{1}=(\lambda_{1,1},\cdots,\lambda_{N,1})^{{}^{\prime}}=\Lambda B_{1}$ with $B_{1}=[1,0,0;0,1,0;0,0,0]$. For the second regime, we set $\Lambda_{2}=(\lambda_{1,2},\cdots,\lambda_{N,2})^{{}^{\prime}}=\Lambda B_{2}$, where $B_{2}=[2,x,y;0,2,z;0,0,0]$ and $x$, $y$, $z$ are i.i.d. $N(0,1)$. For the third regime, we set $\Lambda_{3}=(\lambda_{1,3},\cdots,\lambda_{N,3})^{{}^{\prime}}=\Lambda B_{3}$ with $B_{3}=[0,0,0;0,0,0;0,0,1]$. In this setup, we can find a nonsingular matrix $\bf R$} such that
[…]
The first breakpoint in this setting is a Type 2 change with $r=3>r_{1,2}=r_{1}=r_{2}=2$. The second breakpoint, however, is a Type 1 change and is responsible for the rank deficiency of the pseudo-factors in the first two regimes. As shown in Table 5, our QML estimator yields near-zero estimation errors for the second breakpoint (Type 1) and bounded errors for the first (Type 2), corroborating the theoretical findings in Theorem 2. In contrast, both $\hat{k}_{BKW}$ and $\hat{k}_{MS}$ exhibit substantially larger estimation errors than our $\hat{k}_{QML}$.
Table 5: Simulated root mean squared errors (RMSEs) and mean absolute errors (MAEs) of $\hat{k}_{QML}$, $\hat{k}_{BKW}$ and $\hat{k}_{MS}$, under DGP 1.E.
$N,T$ $(\hat{k}_{1},\hat{k}_{2})_{QML}$ $(\hat{k}_{1},\hat{k}_{2})_{BKW}$ $(\hat{k}_{1},\hat{k}_{2})_{MS}$
RMSE MAE RMSE MAE RMSE MAE
$\rho=0$ $\alpha=0$ $\beta=0$
100,100 (3.633,0.155) (1.949,0.022) (15.145,4.496) (8.667,1.695) (29.270,5.861) (25.831,3.295)
100,300 (2.520,0.148) (1.366,0.022) (8.309,0.780) (3.844,0.300) (48.334,16.411) (36.875,4.296)
300,300 (2.427,0.100) (1.310,0.010) (8.784,1.627) (3.799,0.452) (71.554,4.018) (60.385,0.718)
300,600 (2.174,0.078) (1.225,0.006) (5.545,0.839) (2.842,0.346) (188.222,8.533) (167.173,3.990)
600,600 (2.477,0.084) (1.290,0.007) (8.184,0.884) (3.505,0.366) (212.646,6.323) (207.218,2.655)
$\rho=0.7$ $\alpha=0$ $\beta=0$
100,100 (6.912,0.424) (3.832,0.024) (25.394,9.812) (19.344,5.912) (23.705,6.482) (19.818,3.608)
100,300 (5.773,0.118) (3.134,0.014) (39.502,12.142) (19.971,3.552) (38.722,16.872) (31.729,4.553)
300,300 (6.362,0.071) (3.155,0.005) (39.445,11.937) (19.855,3.580) (91.617,6.655) (79.851,3.000)
300,600 (5.176,0.089) (2.636,0.008) (29.764,9.061) (11.098,1.201) (163.143,7.993) (133.892,3.420)
600,600 (5.373,0.055) (2.730,0.003) (39.830,10.978) (14.478,1.843) (224.301,9.778) (218.798,5.816)
$\rho=0$ $\alpha=0.3$ $\beta=0$
100,100 (3.913,0.675) (1.964,0.058) (15.890,5.301) (9.032,2.079) (26.293,4.959) (22.084,2.821)
100,300 (2.436,0.179) (1.310,0.032) (9.848,2.290) (4.028,0.376) (34.548,15.027) (27.554,3.916)
300,300 (2.653,0.100) (1.367,0.010) (11.124,2.851) (4.060,0.541) (52.027,3.326) (38.841,1.053)
300,600 (2.409,0.100) (1.344,0.010) (5.755,0.925) (3.083,0.364) (67.729,1.197) (55.553,0.678)
600,600 (2.317,0.055) (1.238,0.003) (7.029,0.908) (3.050,0.370) (84.391,2.551) (64.069,0.907)
$\rho=0$ $\alpha=0$ $\beta=0.3$
100,100 (3.280,0.145) (1.713,0.019) (16.181,5.220) (9.266,2.009) (26.854,5.717) (22.728,3.244)
100,300 (2.288,0.145) (1.222,0.019) (9.004,2.638) (3.879,0.435) (41.812,18.103) (31.784,5.114)
300,300 (2.540,0.145) (1.276,0.021) (12.016,3.572) (4.384,0.610) (77.056,2.844) (65.225,0.879)
300,600 (2.093,0.084) (1.141,0.007) (5.856,0.896) (2.963,0.350) (159.473,4.668) (129.832,2.168)
600,600 (2.062,0.071) (1.176,0.005) (6.939,0.985) (3.254,0.399) (211.883,6.943) (203.669,3.291)
$\rho=0.7$ $\alpha=0.3$ $\beta=0.3$
100,100 (7.263,1.070) (3.980,0.083) (25.069,9.428) (19.206,5.684) (21.130,7.205) (17.294,4.071)
100,300 (5.830,0.182) (3.052,0.027) (44.287,14.081) (22.931,4.522) (38.584,18.216) (32.287,5.701)
300,300 (6.004,0.063) (2.937,0.004) (41.843,12.271) (21.719,3.710) (43.257,3.377) (33.133,1.039)
300,600 (5.743,0.045) (2.893,0.002) (34.972,9.347) (13.471,1.366) (72.697,1.046) (59.165,0.464)
600,600 (5.904,0.000) (2.907,0.000) (30.896,9.040) (11.606,1.317) (73.199,2.969) (59.281,0.674)
DGP 2. Table 6 reports the percentage of correct detection of the number of breaks under DGP 1.A-E using the information criterion (30) with $m_{max}=5$. It can be seen that the proposed information criterion can almost correctly identify the number of breakpoints as $N$ and $T$ increases.
@@
DGP 3. Finally, we assess the accuracy of our information criterion (30) in selecting the number of breakpoints, with $m_{max}=5$. We consider scenarios where the true number of breaks, $m_{0}$, ranges from 0 to 4. The factors are generated by (31) and we set $r_{0}=2$. For each regime, the $N\times 2$ factor loading matrix has elements independently drawn from $N(0,1/r_{0})$. This implies that the number of pseudo-factors, $r$, is $(m_{0}+1)r_{0}$.
Before estimating the factors and factor loadings, we use Bai and Ng’s (2002) IC to determine the number of factors for the full samples with $r_{max}=12$. Notably, if no breakpoint exists, the criterion function in (30) is simply zero, as the sample covariance of the factors becomes an identity matrix. As Table 7 illustrates, the probability of our proposed information criterion accurately selecting the true number of breakpoints approaches one as the sample size increases. This finding further validates our Theorem 3.
Table 7: Percentage of correct detection of the number of breaks under different number of breaks.
$N,T$ 0 1 2 3 4
@@ 6 Empirical application
In this section we apply our method to detect breakpoints for the FRED-MD (Federal Reserve Economic Data - Monthly Data) data set. The FRED database is maintained by the Research division of the Federal Reserve Bank of St. Louis, and is publicly accessible and updated in real-time. Our analysis utilizes a sample of 126 unbalanced monthly time series, covering the period from January 1959 to July 2024. These series are categorized into eight groups: (1) output and income, (2) labor market, (3) housing, (4) consumption, orders, and inventories, (5) money and credit, (6) interest rates and exchange rates, (7) prices, and (8) stock market. Following the literature, we replaced the missing values replaced by the EM algorithm and transform the data into approximately stationary series and replaced outliers, so that we obtain a total of $T=785$ monthly observations for each macroeconomic variable. One can refer to McCracken and Ng (2016) for the detailed data description.
@@
We use $h=20$ as the minimum distance between breakpoints. We determine the number of factors for the entire sample using the IC2 criterion of Bai and Ng (2002), which yield seven factors. Subsequently, we apply our proposed information criterion (30) and the QML method to identify the number of breaks and estimate the breakpoint locations. The estimated breakpoints are January 1969, January 1983, June 2008, March 2010, and February 2020. These breakpoints correspond to significant economic events. January 1969 marked a surge in U.S. inflation, primarily driven by increased spending on the Vietnam War and “Great Society” programs. January 1983 roughly coincides with the onset of the Great Moderation, a period characterized by reduced macroeconomic volatility. June 2008 signified the early stages of the Global Financial Crisis, as the collapse of the housing market began to impact the broader financial sector. March 2010 is an approximate to the end of the Great Recession for the U.S. economy. February 2020 represented the initial economic disruptions stemming from the COVID-19 pandemic, which rapidly led to a sharp global downturn.
Finally, based on the five breakpoints, we divided the entire period into six regimes and applied the IC2 criterion from Bai and Ng (2002) to estimate the number of factors in each regime, yielding 2, 5, 7, 3, 4, and 7, respectively. The variation in the number of factors across adjacent regimes indicates that all identified breaks are Type 1 changes.

### v3

@@ 5 Simulation
Table 1: Simulated root mean squared errors (RMSEs) of $\hat{k}_{QML}$, $\hat{k}_{BKW}$, and $\hat{k}_{MS}$ under DGPs 1.A–1.E
QML BKW MS
DGP ($N,T$) $(\hat{k}_{1},\hat{k}_{2})$ $(\hat{k}_{1},\hat{k}_{2})$ $(\hat{k}_{1},\hat{k}_{2})$
DGP 1.A ($b=1$)
100,100 (0.118, 0.095) (2.471, 3.445) (2.121, 5.266)
100,300 (0.055, 0.055) (2.317, 1.338) (0.744, 0.532)
300,300 (0.000, 0.055) (1.167, 1.200) (0.969, 0.693)
300,600 (0.045, 0.032) (1.046, 1.051) (0.985, 0.823)
600,600 (0.032, 0.032) (0.818, 1.030) (1.044, 0.870)
DGP 1.A ($b=0$)
100,100 (0.100, 0.114) (3.115, 2.989) (4.336, 1.265)
100,300 (0.089, 0.141) (1.159, 0.967) (0.933, 0.880)
300,300 (0.000, 0.055) (1.099, 2.137) (1.153, 0.920)
300,600 (0.032, 0.032) (0.921, 1.050) (1.298, 0.943)
600,600 (0.032, 0.000) (0.788, 0.980) (1.429, 0.916)
DGP 1.B
100,100 (0.531, 0.577) (21.127, 19.862) (12.497, 13.936)
100,300 (0.504, 0.427) (30.374, 30.487) (31.692, 31.232)
300,300 (0.344, 0.336) (30.739, 26.776) (23.455, 21.124)
300,600 (0.295, 0.321) (16.831, 17.168) (5.083, 6.411)
600,600 (0.221, 0.307) (15.492, 18.497) (3.111, 3.017)
DGP 1.C
100,100 (1.097, 0.249) (17.995, 21.305) (14.184, 3.582)
100,300 (0.982, 0.278) (35.058, 24.497) (32.378, 24.107)
300,300 (0.602, 0.138) (31.680, 20.713) (33.039, 2.012)
300,600 (0.518, 0.127) (35.115, 8.706) (53.368, 1.914)
600,600 (0.486, 0.123) (31.109, 9.982) (59.029, 2.181)
DGP 1.D
100,100 (8.565, 8.408) (15.741, 16.328) (16.173, 16.614)
100,300 (5.619, 6.389) (28.825, 30.149) (45.309, 44.115)
300,300 (6.326, 5.310) (29.341, 28.247) (41.986, 44.212)
300,600 (4.921, 5.171) (22.760, 22.752) (93.980, 84.635)
600,600 (5.691, 5.362) (19.944, 19.463) (82.489, 86.675)
DGP 1.E
100,100 (7.263, 1.070) (25.069, 9.428) (21.130, 7.205)
100,300 (5.830, 0.182) (44.287, 14.081) (38.584, 18.216)
300,300 (6.004, 0.063) (41.843, 12.271) (43.257, 3.377)
300,600 (5.743, 0.045) (34.972, 9.347) (72.697, 1.046)
600,600 (5.904, 0.000) (30.896, 9.040) (73.199, 2.969)
Finally, we assess the precision of our information criterion (30) in selecting the number of breaks under DGPs 1.A–E and 2 (with $m_{max}=5$). Table 2 reports the percentage of correct detection of the number of breaks under DGP 1.A–E (all featuring two breaks) using the information criterion (30). It can be seen that the proposed information criterion can approximately correctly identify the number of breakpoints as $N$ and $T$ increase.
DGP 2. In addition, DGP 2 considers scenarios where the true number of breaks, $m_{0}$, ranges from 0 to 4. The factors are generated by (31) and we set $r_{0}=2$. For each regime, the $N\times 2$ factor loading matrix has elements independently drawn from $N(0,1/r_{0})$. This implies that the number of pseudo-factors, $r$, is $(m_{0}+1)r_{0}$.
Before estimating the factors and factor loadings, we use Bai and Ng’s (2002) information criterion to determine the number of factors for the full samples with $r_{max}=12$. Notably, if no breakpoint exists, the criterion function in (30) is simply zero, as the sample covariance of the factors becomes an identity matrix. As Table 3 illustrates, the probability of our proposed information criterion accurately selecting the true number of breakpoints approaches one as the sample size increases. This finding further validates our Theorem 3.
@@
Table 3: Percentage of correct detection of the number of breaks for different $m_{0}$.
$N,T$ $m_{0}=$0 $m_{0}=$1 $m_{0}=$2 $m_{0}=$3 $m_{0}=$4
@@ 6 Empirical application
In this section we apply our method to detect breakpoints for the FRED-MD (Federal Reserve Economic Data - Monthly Data) data set. The FRED database is maintained by the research division of the Federal Reserve Bank of St. Louis, and is publicly accessible and updated in real time. Our analysis utilizes a sample of 126 unbalanced monthly time series, covering the period from January 1959 to July 2024. These series are categorized into eight groups: (1) output and income, (2) labor market, (3) housing, (4) consumption, orders, and inventories, (5) money and credit, (6) interest rates and exchange rates, (7) prices, and (8) stock market. Following the literature, we replace the missing values using the expectation-maximization algorithm and transform the data into approximately stationary series with the outliers replaced, so that we obtain a total of $T=785$ monthly observations for each macroeconomic variable. One can refer to McCracken and Ng (2016) for the detailed data description.
In selecting the number of breaks, we set the minimum segment length to $h=20$ months and restrict candidate break dates to the monthly grid. We consider $m\in\{0,1,\ldots,8\}$ as the candidate numbers of breaks. For each $m$, we minimize the QML objective over all admissible partitions using the dynamic programming algorithm and then select $\hat{m}$ by minimizing Eq. (30). In our empirical application, $N=126$, $r=7$ (selected by Bai–Ng IC2 on the full sample), and $\hat{\rho}=0.9574$, so the penalty term is numerically
$\log(\min\{N,T\})\,r^{2}(1+|\hat{\rho}|)=\log(126)\times 7^{2}\times(1+0.9574)\approx 463.86.$
@@
Hence the penalty used in the information criterion is $463.86\times m$. No additional tuning constant is used. The selected model yields five breakpoints: January 1969, January 1983, June 2008, March 2010, and February 2020. These breakpoints correspond to significant economic events. January 1969 marked a surge in U.S. inflation, primarily driven by increased spending on the Vietnam War and the “Great Society” programs. January 1983 roughly coincides with the onset of the Great Moderation, a period characterized by reduced macroeconomic volatility. June 2008 signifies the early stages of the Global Financial Crisis, as the collapse of the housing market began to impact the broader financial sector. March 2010 is approximately the end of the Great Recession for the U.S. economy. February 2020 represents the initial economic disruptions stemming from the COVID-19 pandemic, which rapidly led to a sharp global downturn.
Using the estimated breakpoints, we split the sample into six regimes and apply Bai–Ng IC2 within each regime, obtaining 2, 5, 7, 3, 4, and 7 factors, respectively. Under our classification rule, the variation in the estimated numbers of factors across adjacent regimes suggests that all identified breaks are Type 1 changes. For clarity, the regime with three factors is 2008:07–2010:03, whereas the following regime 2010:04–2020:02 contains four factors. Since principal components are identified only up to rotation, the economic meaning of individual factors is not always clear, and higher-order components are particularly difficult to interpret. We therefore focus on a more stable feature of the data: the concentration of explained variance in the leading principal components.
This concentration rises substantially in the crisis regimes. In 2008:07–2010:03, the first principal component explains about $34.7\%$ of total variance and the first three explain about $59.4\%$. In 2020:03–2024:07, the first principal component alone explains about $52.6\%$ of total variance. By contrast, in the non-crisis regimes 1983:02–2008:06 and 2010:04–2020:02, the first principal component explains only about $10.0\%$ and $13.2\%$, respectively. This indicates that crisis periods are dominated by a small number of strong common shocks, whereas common variation is more diffuse in normal times.
The two crisis episodes also differ in the composition of the leading common variation. During 2008:07–2010:03, the leading common variation is more closely associated with financial and price-related categories, including money and credit, interest rates and exchange rates, and prices. During 2020:03–2024:07, the leading common variation is much more strongly concentrated in output and income, as well as labor-market variables. This difference is consistent with the distinct nature of the Global Financial Crisis and the COVID-19 shock. Detailed regime-wise explained-variance results are reported in Appendix E of the supplementary material.